A medicine capsule is in the shape of a cylinder with two hemispheres struck to each of its end. The length of entire capsule is 14 mm and the diameter of the capsule is 5 mm. Find the surface area
Answer by mananth(16946)

Height of cylinder = 14-5 = 9 mm
Radius = 2.5 mm
Curved surface area = 2*pi*2.5*9= 45pi
curved surface area of two hemispheres = 2pir^2*2
=4*pi*2.5^2*
=25pi
Total curved surface area of capsule = 45pi+25pi=70 pi mm^2
220mm^2
m.ananth@hotmail.ca